OpenStudy (highschoolmom2010):

A sphere has the volume given. Find its surface area to the nearest whole number. V=900 in ^3

OpenStudy (kagıtucak):

V=4/3*pi*r^3 S=4*pi*r^2 you have to use these formulas

OpenStudy (highschoolmom2010):

i dont know how to use volume to find surface area

OpenStudy (anonymous):

The surface area's formula is 4 pi r 2. The r is the radius. You can simply plug the numbers into the equation and round it to get your answer.

OpenStudy (primeralph):